# SwitchLoRA: Switched Low-Rank Adaptation Can Learn Full-Rank Information

### v2

@@ Abstract
In the training of large language models, parameter-efficient techniques such as LoRA optimize memory usage and reduce communication overhead during the fine-tuning phase. However, applying such techniques directly during the pre-training phase results in poor performance, primarily because the premature implementation of low-rank training significantly reduces model accuracy. Existing methods like ReLoRA and GaLore have attempted to address this challenge by updating the low-rank subspace. However, they still fall short of achieving the accuracy of full-rank training because they must limit the update frequency to maintain optimizer state consistency, hindering their ability to closely approximate full-rank training behavior. In this paper, we introduce SwitchLoRA, a parameter-efficient training technique that frequently and smoothly replaces the trainable parameters of LoRA adapters with alternative parameters. SwitchLoRA updates the low-rank subspace incrementally, targeting only a few dimensions at a time to minimize the impact on optimizer states. This allows a higher update frequency, thereby enhancing accuracy by enabling the updated parameters to more closely mimic full-rank behavior during the pre-training phase. Our results demonstrate that SwitchLoRA actually surpasses full-rank training, reducing perplexity from 15.23 to 15.01 on the LLaMA 1.3B model while reducing communication overhead by 54% on the LLaMA 1.3B model. Furthermore, after full fine-tuning the SwitchLoRA pre-trained model and the full-rank pre-trained model on the GLUE benchmark, the SwitchLoRA pre-trained model showed an average accuracy gain of about 1% over the full-rank pre-trained model. This demonstrates enhanced generalization and reasoning capabilities of SwitchLoRA.
@@ 1 Introduction
The size of large language models (LLMs) has increased rapidly due to the advent of the transformer architecture [47]. To support the training of large models, distributed training techniques such as data parallelism [5, 29], tensor parallelism [41], pipeline parallelism [22, 39] and the Zero Redundancy Optimizer [40] have been employed. However, distributed training of trillion-scale models incurs significant inter-node communication overhead from synchronizing extensive parameter gradients across multiple nodes.
@@
To address these challenges, various parameter-efficient strategies have been proposed. Techniques such as model sparsification [1, 42] and progressive model pruning during training [11] have shown promise. Additionally, methods leveraging Singular Value Decomposition (SVD) to approximate full-rank matrices in low-rank spaces have been explored [43, 49, 59]. Beyond the entire training process, several techniques improve adaptability and efficiency during the fine-tuning phase. For example, methods such as the Adapter [20, 18] and Prefix-tuning [30] introduce additional trainable layers while freezing the remaining parameters.
@@
Another noteworthy fine-tuning strategy is Low-Rank Adaptation (LoRA) [21], which introduces no computational overhead during inference while maintaining training accuracy. However, previous studies [49, 50, 34] have observed that parameter-efficient methods such as LoRA perform less efficiently during the pre-training phase because the premature use of low-rank training leads to a considerable loss in model accuracy. To increase the rank of updated parameters and benefit from low-rank training, ReLoRA [34] applies the structure of LoRA and periodically resets LoRA adapters. Similarly, GaLore [60] projects gradients onto a subspace, updating this subspace periodically. These approaches update the descent direction of trainable parameters to mimic the behavior of full-rank training, thereby overcoming the limitations observed in existing implementations of low-rank adaptation. However, we find that the intervals between resetting/updating steps in ReLoRA and GaLore are set to relatively large values because too frequent changes in the updating direction can cause inconsistency in optimizer states, which may not sufficiently approximate the behavior of full-rank training, resulting in a loss of accuracy.
@@
To address this challenge, as illustrated in Figure 1, we introduce SwitchLoRA, which enables smooth and frequent adjustments to the trainable parameters of the LoRA matrices while introducing negligible additional computational overhead. SwitchLoRA maintains a set of candidate vectors for each matrix within the LoRA adapters. At each training step, it replaces portions of the column or row vectors with these candidate vectors, subsequently training the LoRA adapters. This process minimizes the impact on optimizer states, thus allowing for a higher update frequency compared to ReLoRA and GaLore. By more closely approximating full-rank parameter updating behaviors during the pre-training phase, this approach enhances overall accuracy.
@@
  • We propose SwitchLoRA to facilitate smooth and frequent adjustments to the trainable parameters of the LoRA matrices through low-rank adaptation, maintaining the accuracy of full-rank training while reducing communication overhead.
@@
  • We experimentally validate SwitchLoRA on various sizes of the LLaMA model. SwitchLoRA shows significant perplexity improvements when compared to ReLoRA [34] and GaLore [60]. For the 1.3B model, SwitchLoRA achieves a perplexity of 15.01, surpassing the 15.23 perplexity obtained with full-rank training. Furthermore, by performing full fine-tuning on the resulting 1.3B model using the GLUE [48] tasks to validate the reasoning capabilities, we demonstrate that SwitchLoRA enhances model accuracy by approximately 1% on average, compared to the full-rank training method.
@@ 2 Methodology
A substantial body of research, such as various pruning methods [15, 4], has demonstrated that neural networks tend to exhibit low-rank characteristics after certain stages of training. Techniques for parameter-efficient fine-tuning, such as LoRA, capitalize on this observation. Concurrently, studies like [33, 13] have revealed that overparameterization in neural networks can lead to implicit regularization, thereby enhancing generalization. These findings underscore the importance of training with full parameters during the initial phase. Further empirical evidence supporting this phenomenon is provided in works like [49, 50, 34, 60]. Based on these insights, this section proposes a method designed to train a substantial number of parameters while selectively updating only a portion of the parameters at any one time to reduce communication overhead.
@@ 2.1 Low-Rank Adaptation (LoRA)
Introduced in [21], LoRA is designed specifically for the fine-tuning stage of model training.
Consider a pre-trained model with a weight matrix $\mathbf{W}\in\mathbb{R}^{m\times n}$ from a specific linear layer. LoRA proposes an innovative modification: transforming $\mathbf{W}$ into $\mathbf{W}+\frac{\alpha}{r}\mathbf{B}\mathbf{A}$. Here, $\mathbf{B}\in\mathbb{R}^{m\times r}$ and $\mathbf{A}\in\mathbb{R}^{r\times n}$ are newly introduced matrices, where $r$ is a positive integer significantly smaller than both $m$ and $n$. And $\alpha$ is a constant hyperparameter, set to $r$ in the following description to clarify the algorithm’s mechanics. Then during fine-tuning, $\mathbf{W}$ is kept frozen while matrices $\mathbf{B}$ and $\mathbf{A}$ are trained. At the inference stage, $\mathbf{B}\mathbf{A}$ is added to $\mathbf{W}$ which preserves the model’s original structure. The matrix $\mathbf{A}$ is initialized using Kaiming initialization [17], while $\mathbf{B}$ is initially set to a zero matrix to ensure consistency.
@@ 2.2 SwitchLoRA
Below, we detail our proposed SwitchLoRA algorithm, the steps of which are outlined in Algorithm 1 and Algorithm 2.
@@
If we adopt the strategy in LoRA to add $\mathbf{B}\mathbf{A}$ to $\mathbf{W}$ and only update $\mathbf{B}$ and $\mathbf{A}$ from the pre-training stage, according to (2), the rank of updated parameters of the local linear system through the entire training process will be limited to $2r$. This limitation can potentially impede the training efficacy. To mitigate this issue, we alter the values of $\mathbf{b}_{k}$ and $\mathbf{a}_{k}$ to $\mathbf{b}^{\prime}_{k}\in\mathcal{C}(\mathbf{B})$ and $\mathbf{a}^{\prime}_{k}\in\mathcal{C}(\mathbf{A}^{T})$ at appropriate frequencies, respectively, with these new values randomly selected from predefined candidate vectors list $\mathcal{C}(\mathbf{B})$ and $\mathcal{C}(\mathbf{A}^{T})$(one of $\mathbf{b}_{k}^{\prime}$ or $\mathbf{a}_{k}^{\prime}$ can be the same as $\mathbf{b}_{k}$ or $\mathbf{a}_{k}$). To maintain the consistency of the model’s output, we adjust $\mathbf{W}$ by adding the difference between the old and new LoRA components. To be more precise, when $\mathbf{b}_{k}$ and $\mathbf{a}_{k}$ are updated to $\mathbf{b}^{\prime}_{k}$ and $\mathbf{a}^{\prime}_{k}$, we accordingly adjust $\mathbf{W}$ with the equation $\mathbf{W}\leftarrow\mathbf{W}+\mathbf{b}_{k}\mathbf{a}_{k}^{T}-\mathbf{b}_{k}^{\prime}\mathbf{a}_{k}^{\prime T}$.
When implementing these updates, the updated parameters of both $\mathbf{B}$ and $\mathbf{A}$ are derived from $\min(m,n)$ distinct candidate vectors, which ensures updated parameters are full-rank. Readers can refer to [34, 62, 54] for more details.
When selecting candidate vectors, we have the option to choose randomly from $C(\mathbf{B})$ or $C(\mathbf{A}^{T})$. Alternatively, we can select candidate vectors sequentially from $C(\mathbf{B})$ or $C(\mathbf{A}^{T})$, restarting from the beginning once the end of the set is reached. We find that varying the matching orders of vectors $\mathbf{b}_{k}$ and $\mathbf{a}_{k}$ yields only minor differences in outcomes. A theoretical explanation for this phenomenon is provided in Appendix A. Additionally, to conserve GPU memory, spare candidate vectors can be offloaded to the CPU.
@@
As mentioned in [11, 49, 34], the model initially exhibits full internal rank during pre-training, and the internal rank of each layer decreases progressively over time. Consequently, we have adopted an exponential decay function for the switching frequency, namely $frequency=Ce^{-\theta step}$, where the coefficients are determined empirically. Besides, the selection of LoRA rank $r$ for $\mathbf{B}\mathbf{A}$ is influenced by the final internal rank of the layers, which has been extensively explored in [21, 46, 57].
@@
Currently Large Language Models (LLMs) predominantly utilize Adam [26] and AdamW [36] optimizers over SGD, which rely on optimizer states. It is crucial to note that after switching LoRA vectors, the gradients associated with these parameters are also changed, which prevents the reuse of optimizer states. To address this issue, when $\mathbf{a}_{k}$ is switched, we reset the optimizer states of $\mathbf{b}_{k}$. And conversely, when $\mathbf{b}_{k}$ is switched, we reset optimizer states of $\mathbf{a}_{k}$. Note that we reset optimizer states of counterpart pair rather than optimizer states of the switched parameters itself. This approach will be further explained in Appendix A. Additionally, when the optimizer states are reset to zero, we freeze corresponding parameters for $N$ steps to maintain the robustness of the training. In this study, $N$ is set to $5$.
Algorithm 1 Switch algorithm: $\mathbf{W},\mathbf{P},\mathbf{Q}=\text{switch}(\mathbf{W},\mathbf{P},\mathbf{Q},i,j)$. $\mathcal{C}(\mathbf{P})[i]$ is $i$-th predefined candidate vectors for $\mathbf{P}$.
@@
Results in [16, 56] have demonstrated the importance of initialization of LoRA matrices $\mathbf{B}$ and $\mathbf{A}$ to the training effects. Unlike these works, which are applied only during the fine-tuning stage, our method is utilized throughout the entire training process. To achieve appropriate initialization for matrices $\mathbf{B}$ and $\mathbf{A}$ along with their candidate vectors, we follow the idea of Xavier initialization [12] and Kaiming initialization [17]. The derived initialization values are presented as follows:
@@ 3 Related work
Numerous studies [6, 44, 53, 23] have demonstrated the effectiveness of using low-rank factorization to approximate the weights of linear layers in deep neural networks. They employ methods such as SVD to achieve a factorization $\mathbf{U}\mathbf{V}$ that minimizes $\|\mathbf{W}-\mathbf{U}\mathbf{V}\|$. Later on, Pufferfish [49] and subsequent work in Cuttlefish [50] employ full-rank training prior to low-rank training to enhance efficiency. Additionally, they introduce adaptive strategies to determine the necessary duration of full-rank training and to select the appropriate rank for each linear layer for SVD. Further developments in this field include InRank [59], which proposes a low-rank training approach based on greedy low-rank learning [32]. Additional research such as [43] integrates orthogonality into the low-rank models to enhance training accuracy, while [19] introduces low-rank ordered decomposition, a generalization of SVD aimed at improving low-rank training efficiency.
@@
After the introduction of LoRA in [21], which facilitated fine-tuning with very few trainable parameters, numerous works are proposed to improve the performance of LoRA. Improvements include better initialization strategies for LoRA matrices as demonstrated in [51, 52, 38]. Additionally, [16, 25] have adjusted learning rates for $\mathbf{B}$ and $\mathbf{A}$ to optimize training outcomes. Other research efforts, such as those in [27, 35], have modified the training process of LoRA. Moreover, some studies, such as [14, 58], focus on training models from scratch within a sparse model structure.
Similar to our approach, various LoRA variants employ strategies to increase the rank of updated parameters by merging parameters of adapters into $\mathbf{W}$. For instance, Chain of LoRA [54] and ReLoRA [34] merge $\mathbf{B}\mathbf{A}$ into $\mathbf{W}$ and restart training at regular intervals. ReLoRA enables low-rank training during the early phases, yet it still requires 33% of the steps to be full-rank training. Delta-LoRA [62], another variant, targets the fine-tuning phase by updating the matrix $\mathbf{W}$ using the gradients from the LoRA matrices $\mathbf{B}$ and $\mathbf{A}$ as they are updated, enhancing accuracy for fine-tuning.
@@
In addition to previously discussed techniques, there are many other methods to compress models during training. For instance, several studies have introduced quantization to LoRA [8, 31, 24], effectively reducing memory overhead during fine-tuning. Other research employs iterative pruning and growth techniques during training [11, 55, 37, 10]. Additionally, some works focus on compressing gradients through quantization [7, 28] or gradient projection [60]. Notably, [60] presents a recent method for training from scratch that utilizes SVD to project gradients into a periodically updated subspace. This approach also enables the addition of quantization, offering enhanced memory efficiency compared to LoRA.

### v3

@@ Abstract
In the training of large language models, parameter-efficient techniques such as LoRA optimize memory usage and reduce communication overhead and memory usage during the fine-tuning phase. However, applying such techniques directly during the pre-training phase results in poor performance, primarily because the premature implementation of low-rank training significantly reduces model accuracy. Existing methods like ReLoRA and GaLore have attempted to address this challenge by updating the low-rank subspace. However, they still fall short of achieving the accuracy of full-rank training. Specifically, ReLoRA restricts the frequency of updates to preserve optimizer states consistency, hindering its ability to closely approximate full-rank training behavior. Meanwhile, GaLore relies on Singular Value Decomposition (SVD) to approximate the full-rank space, which introduces accuracy loss during the approximation process. In this paper, we introduce SwitchLoRA, a parameter-efficient training technique that frequently and smoothly replaces the trainable parameters of LoRA adapters with alternative parameters. SwitchLoRA updates the low-rank subspace incrementally, targeting only a few dimensions at a time to minimize the impact on optimizer states. This allows a higher update frequency, thereby enhancing accuracy by enabling the updated parameters to more closely mimic full-rank behavior during the pre-training phase. Our results demonstrate that SwitchLoRA actually surpasses full-rank training, reducing perplexity from 15.23 to 15.01 on the LLaMA 1.3B model, while also cutting communication overhead by 54% and memory usage by 13%. Furthermore, after full fine-tuning the SwitchLoRA pre-trained model and the full-rank pre-trained model on the GLUE benchmark, the SwitchLoRA pre-trained model showed an average accuracy gain of about 1% over the full-rank pre-trained model. This demonstrates enhanced generalization and reasoning capabilities of SwitchLoRA.
@@ 1 Introduction
The size of large language models (LLMs) has increased rapidly due to the advent of the transformer architecture (Vaswani et al., 2017). To support the training of large models, distributed training techniques such as data parallelism (Dean et al., 2012; Li et al., 2014), tensor parallelism (Shoeybi et al., 2019), pipeline parallelism (Huang et al., 2019; Narayanan et al., 2021) and the Zero Redundancy Optimizer (Rajbhandari et al., 2020) have been employed. However, distributed training of trillion-scale models incurs significant inter-node communication overhead from synchronizing extensive parameter gradients across multiple nodes.
@@
To address these challenges, various parameter-efficient strategies have been proposed. Techniques such as model sparsification (Alistarh et al., 2018; Stich et al., 2018) and progressive model pruning during training (Frankle and Carbin, 2019) have shown promise. Additionally, methods leveraging Singular Value Decomposition (SVD) to approximate full-rank matrices in low-rank spaces have been explored (Sui et al., 2024; Wang et al., 2021; Zhao et al., 2023). Beyond the entire training process, several techniques improve adaptability and efficiency during the fine-tuning phase. For example, methods such as the Adapter (Houlsby et al., 2019; He et al., 2022) and Prefix-tuning (Li and Liang, 2021) introduce additional trainable layers while freezing the remaining parameters.
@@
Another noteworthy fine-tuning strategy is Low-Rank Adaptation (LoRA) (Hu et al., 2022), which introduces no computational overhead during inference while maintaining training accuracy. However, previous studies Wang et al. (2021, 2023); Lialin et al. (2023) have observed that parameter-efficient methods such as LoRA perform less efficiently during the pre-training phase because the premature use of low-rank training leads to a considerable loss in model accuracy. To increase the rank of updated parameters and benefit from low-rank training, ReLoRA (Lialin et al., 2023) applies the structure of LoRA and periodically resets LoRA adapters. These approaches update the gradient descent direction of trainable parameters to mimic the behavior of full-rank training, thereby overcoming the limitations observed in existing implementations of low-rank adaptation. However, we find that the intervals between resetting/updating steps in ReLoRA are set to relatively large values because too frequent changes in the updating direction can cause inconsistency in optimizer states, which can not sufficiently approximate the behavior of full-rank training, resulting in a loss of accuracy. To make the rank of the updated subspace dominant in the whole space, another work GaLore (Zhao et al., 2024b) addresses this by periodically projecting gradients onto a subspace using SVD to make the rank of the updated subspace dominant within the overall space. However, the compression of gradients in this process also leads to accuracy loss.
@@
To address this challenge, as illustrated in Figure 1, we introduce SwitchLoRA, which enables smooth and frequent adjustments to the trainable parameters of the LoRA matrices while introducing negligible additional computational overhead. SwitchLoRA maintains a set of candidate vectors for each matrix within the LoRA adapters. At each training step, it replaces portions of the column or row vectors with these candidate vectors, subsequently training the LoRA adapters. This process minimizes the impact on optimizer states, thus allowing for a higher update frequency compared to ReLoRA. By more closely approximating full-rank parameter updating behaviors during the pre-training phase, this approach enhances overall accuracy.
@@
  • We propose SwitchLoRA to facilitate smooth and frequent adjustments to the trainable parameters of the LoRA matrices through low-rank adaptation, maintaining the accuracy of full-rank training while reducing memory usage and communication overhead.
@@
  • We compare the training time and memory usage of full-rank training, SwitchLoRA, and LoRA. The results indicate that SwitchLoRA and LoRA, when using the same LoRA rank, exhibit nearly identical memory overhead and training time. As the model size increases from 1.3B to 7B, SwitchLoRA reduces memory usage by 13% to 65% compared to full-rank training.
  • We experimentally validate SwitchLoRA on various sizes of the LLaMA model. SwitchLoRA shows significant perplexity improvements when compared to ReLoRA (Lialin et al., 2023) and GaLore (Zhao et al., 2024b). For the 1.3B model, SwitchLoRA achieves a perplexity of 15.01, surpassing the 15.23 perplexity obtained with full-rank training. Furthermore, by performing full fine-tuning on the resulting 1.3B model using the GLUE (Wang et al., 2019) tasks to validate the reasoning capabilities, we demonstrate that SwitchLoRA enhances model accuracy by approximately 1% on average, compared to the full-rank training method.
@@ 2 Methodology
A substantial body of research, such as various pruning methods (Han et al., 2015; Blalock et al., 2020), has demonstrated that neural networks tend to exhibit low-rank characteristics after certain stages of training. Techniques for parameter-efficient fine-tuning, such as LoRA, capitalize on this observation. Concurrently, studies like Li et al. (2020); Gunasekar et al. (2017) have revealed that overparameterization in neural networks can lead to implicit regularization, thereby enhancing generalization. These findings underscore the importance of training with full parameters during the initial phase. Further empirical evidence supporting this phenomenon is provided in works like Wang et al. (2021, 2023); Lialin et al. (2023); Zhao et al. (2024b). Based on these insights, this section proposes a method designed to train a substantial number of parameters while selectively updating only a portion of the parameters at any one time to reduce memory usage and communication overhead.
@@ 2.1 Low-Rank Adaptation (LoRA)
Introduced in Hu et al. (2022), LoRA is designed specifically for the fine-tuning stage of model training.
Consider a pre-trained model with a weight matrix $\mathbf{W}\in\mathbb{R}^{m\times n}$ from a specific linear layer. LoRA proposes an innovative modification: transforming $\mathbf{W}$ into $\mathbf{W}+\frac{\alpha}{r}\mathbf{B}\mathbf{A}$. Here, $\mathbf{B}\in\mathbb{R}^{m\times r}$ and $\mathbf{A}\in\mathbb{R}^{r\times n}$ are newly introduced matrices, where $r$ is a positive integer significantly smaller than both $m$ and $n$. And $\alpha$ is a constant hyperparameter, set to $r$ in the following description to clarify the algorithm’s mechanics. The matrix $\mathbf{A}$ is initialized using Kaiming initialization (He et al., 2015), while $\mathbf{B}$ is initially set to a zero matrix to ensure consistency. During fine-tuning, $\mathbf{W}$ is kept frozen while matrices $\mathbf{B}$ and $\mathbf{A}$ are trained. At the inference stage, $\frac{\alpha}{r}\mathbf{B}\mathbf{A}$ is added to $\mathbf{W}$ which preserves the model’s original structure.
@@ 2.2 SwitchLoRA
The training process of SwitchLoRA is identical to that of LoRA, except that SwitchLoRA switches between LoRA vectors and candidate vectors at each training step. Below, we detail our proposed SwitchLoRA algorithm, the steps of which are outlined in Algorithm 1 and Algorithm 2.
@@
If we adopt the strategy in LoRA to add $\mathbf{B}\mathbf{A}$ to $\mathbf{W}$ and only update $\mathbf{B}$ and $\mathbf{A}$ from the pre-training stage, according to (2), the rank of updated parameters of the local linear system through the entire training process will be limited to $2r$. This limitation can potentially impede the training efficacy. To mitigate this issue, we alter the values of $\mathbf{b}_{k}$ and $\mathbf{a}_{k}$ to $\mathbf{b}^{\prime}_{k}\in\mathcal{C}(\mathbf{B})$ and $\mathbf{a}^{\prime}_{k}\in\mathcal{C}(\mathbf{A}^{T})$ at appropriate frequencies, respectively, with these new values selected from predefined candidate vectors list $\mathcal{C}(\mathbf{B})$ and $\mathcal{C}(\mathbf{A}^{T})$(one of $\mathbf{b}_{k}^{\prime}$ or $\mathbf{a}_{k}^{\prime}$ can be the same as $\mathbf{b}_{k}$ or $\mathbf{a}_{k}$). To maintain the consistency of the model’s output, we adjust $\mathbf{W}$ by adding the difference between the old and new LoRA components. To be more precise, when $\mathbf{b}_{k}$ and $\mathbf{a}_{k}$ are updated to $\mathbf{b}^{\prime}_{k}$ and $\mathbf{a}^{\prime}_{k}$, we accordingly adjust $\mathbf{W}$ with the equation $\mathbf{W}\leftarrow\mathbf{W}+\mathbf{b}_{k}\mathbf{a}_{k}^{T}-\mathbf{b}_{k}^{\prime}\mathbf{a}_{k}^{\prime T}$.
When implementing these updates, the updated parameters of both $\mathbf{B}$ and $\mathbf{A}$ are derived from $\min(m,n)$ distinct candidate vectors, which ensures updated parameters are full-rank. Readers can refer to Lialin et al. (2023); Zi et al. (2023); Xia et al. (2024) for more details.
When selecting candidate vectors, we have the option to choose randomly from $C(\mathbf{B})$ or $C(\mathbf{A}^{T})$. Alternatively, we can select candidate vectors sequentially from $C(\mathbf{B})$ or $C(\mathbf{A}^{T})$, restarting from the beginning once the end of the set is reached. We find that varying the matching orders of vectors $\mathbf{b}_{k}$ and $\mathbf{a}_{k}$ yields only minor differences in outcomes. A theoretical explanation for this phenomenon is provided in Appendix A. Additionally, to conserve GPU memory, spare candidate vectors are offloaded to the CPU.
@@
As mentioned in Frankle and Carbin (2019); Wang et al. (2021); Lialin et al. (2023), the model initially exhibits full internal rank during pre-training, and the internal rank of each layer decreases progressively over time. Consequently, we have adopted an exponential decay function for the switching frequency, namely $frequency=Ce^{-\theta step}$, where the coefficients are determined empirically. Besides, the selection of LoRA rank $r$ for $\mathbf{B}\mathbf{A}$ is influenced by the final internal rank of the layers, which has been extensively explored in Hu et al. (2022); Valipour et al. (2023); Zhang et al. (2023b).
@@
Currently Large Language Models (LLMs) predominantly utilize Adam (Kingma and Ba, 2015) and AdamW (Loshchilov and Hutter, 2019) optimizers over SGD, which rely on optimizer states. It is crucial to note that after switching LoRA vectors, the gradients associated with these parameters are also changed, which prevents the reuse of optimizer states. To address this issue, when $\mathbf{a}_{k}$ is switched, we reset the optimizer states of $\mathbf{b}_{k}$. And conversely, when $\mathbf{b}_{k}$ is switched, we reset optimizer states of $\mathbf{a}_{k}$. Note that we reset optimizer states of counterpart pair rather than optimizer states of the switched parameters itself. This approach will be further explained in Appendix A. Additionally, when the optimizer states are reset to zero, we freeze corresponding parameters for $N$ steps to maintain the robustness of the training. In this study, $N$ is set to $5$.
Algorithm 1 Switching algorithm: $\mathbf{W},\mathbf{P},\mathbf{Q}=\text{switch}(\mathbf{W},\mathbf{P},\mathbf{Q},i,j)$. $\mathcal{C}(\mathbf{P})[i]$ is $i$-th predefined candidate vectors for $\mathbf{P}$. All candidate vectors in $\mathcal{C}(\mathbf{P})$ are stored in the CPU and transferred to the GPU as needed.
@@
Results in Hayou et al. (2024); Zhang et al. (2023a) have demonstrated the importance of initialization of LoRA matrices $\mathbf{B}$ and $\mathbf{A}$ to the training effects. Unlike these works, which are applied only during the fine-tuning stage, our method is utilized throughout the entire training process. To achieve appropriate initialization for matrices $\mathbf{B}$ and $\mathbf{A}$ along with their candidate vectors, we follow the idea of Xavier initialization (Glorot and Bengio, 2010) and Kaiming initialization (He et al., 2015). Specifically, the values of $\mathbf{B}$ and $\mathbf{A}$ are randomly initialized using a uniform distribution with zero mean and the following standard variance:
@@ 3 Related work
Numerous studies (Denton et al., 2014; Tai et al., 2016; Wen et al., 2017; Idelbayev and Carreira-Perpinán, 2020) have demonstrated the effectiveness of using low-rank factorization to approximate the weights of linear layers in deep neural networks. They employ methods such as SVD to achieve a factorization $\mathbf{U}\mathbf{V}$ that minimizes $\|\mathbf{W}-\mathbf{U}\mathbf{V}\|$. Later on, Pufferfish (Wang et al., 2021) and subsequent work in Cuttlefish (Wang et al., 2023) employ full-rank training prior to low-rank training to enhance efficiency. Additionally, they introduce adaptive strategies to determine the necessary duration of full-rank training and to select the appropriate rank for each linear layer for SVD. Further developments in this field include InRank (Zhao et al., 2023), which proposes a low-rank training approach based on greedy low-rank learning (Li et al., 2021). Additional research such as Sui et al. (2024) integrates orthogonality into the low-rank models to enhance training accuracy, while Horváth et al. (2024) introduces low-rank ordered decomposition, a generalization of SVD aimed at improving low-rank training efficiency.
@@
After the introduction of LoRA in Hu et al. (2022), which facilitated fine-tuning with very few trainable parameters, numerous works are proposed to improve the performance of LoRA. Improvements include better initialization strategies for LoRA matrices as demonstrated in Wang et al. (2024); Wang and Liang (2024); Meng et al. (2024). Additionally, Hayou et al. (2024); Kalajdzievski (2023) have adjusted learning rates for $\mathbf{B}$ and $\mathbf{A}$ to optimize training outcomes. Other research efforts, such as those in Kopiczko et al. (2024); Liu et al. (2024), have modified the training process of LoRA. Moreover, some studies, such as Han et al. (2024); Zhao et al. (2024a), focus on training models from scratch within a sparse model structure.
Similar to our approach, various LoRA variants employ strategies to increase the rank of updated parameters by merging parameters of adapters into $\mathbf{W}$. For instance, Chain of LoRA (Xia et al., 2024) and ReLoRA (Lialin et al., 2023) merge $\mathbf{B}\mathbf{A}$ into $\mathbf{W}$ and restart training at regular intervals. ReLoRA enables low-rank training during the early phases, yet it still requires 33% of the steps to be full-rank training to maintain model accuracy. Delta-LoRA (Zi et al., 2023), another variant, targets the fine-tuning phase by updating the matrix $\mathbf{W}$ using the gradients from the LoRA matrices $\mathbf{B}$ and $\mathbf{A}$ as they are updated, enhancing accuracy for fine-tuning.
@@
In addition to previously discussed techniques, there are many other methods to compress models during training. For instance, several studies have introduced quantization to LoRA (Dettmers et al., 2023; Li et al., 2023b; Jeon et al., 2024), effectively reducing memory overhead during fine-tuning. Other research employs iterative pruning and growth techniques during training (Frankle and Carbin, 2019; You et al., 2019; Lym et al., 2019; Evci et al., 2020). Additionally, some works focus on compressing gradients through quantization (Dettmers et al., 2022; Li et al., 2023a) or gradient projection (Zhao et al., 2024b). Notably, Zhao et al. (2024b) presents a recent method for training from scratch that utilizes SVD to project gradients into a periodically updated subspace. This approach also enables the addition of quantization, offering enhanced memory efficiency compared to LoRA.